OpenStudy (anonymous):

@whpalmer4 half life of element is about 1200 years. after 3700 years what percentage P of a sample of this element remains? (round to nearest 10th)

OpenStudy (whpalmer4):

Okay, we're halving the amount every 1200 years, so you can just write down the equation: \[P(t) = 2^{-t/1200}\] At \(t=0\), that gives you \(P(t) = 2^0 = 1\). At \(t=1200\), that gives you \(P(1200) = 2^{-1200/1200} = 2^{-1} = 0.5\) At \(t = 2400\), that gives you \(P(2400)=2^{-2400/1200} = 2^{-2} = 0.25\) All you have to do is compute \[P(3700) = 2^{-3700/1200} =\]

OpenStudy (whpalmer4):

As \(3600/1200 = 3\), this should be just a bit less than \(2^{-3} = 0.125\) Don't forget to multiply your decimal by 100% to get a percentage, and round to the appropriate place.

OpenStudy (anonymous):

so .1179842891 rounded to tenth is just .1 then?

OpenStudy (whpalmer4):

0.117984*100% = 11.7984%, round that to the nearest 10th...